5 years ago, Barton Industries issued 25-year noncallable, semiannual bonds with a $1,000 face value and a 9% coupon, semiannual payment ($45 payment every 6 months). The bonds currently sell for $844.87. If the firm's marginal tax rate is 25%, what is the firm's after-tax cost of debt? Do not round intermediate calculations. Round your answer to two decimal places.

  %
Answer #1

using excel rate function

Yield to maturity  

= RATE(number_of_periods, payment_per_period, present_value, [future_value], [end_or_beginning], [rate_guess]) * compounds in aa year

= RATE(50,45,-844.87,1000) * 2

= 10.81%

after tax cost of debt = 10.81% * (1-0.25)

= 8.1075%

= 8.11%
Answer #2

SOLUTION :


Using online calculator for YTM ,  ref : https://dqydj.com/bond-yield-to-maturity-calculator/ :


Key -in the following :


Current market price ($) = 844.87


Bond Face Value ($) = 1000


Years to maturity (25 - 5) = 20 


Annual coupon rate (%) = 9.0


Frequency of coupon = 2 (being semi-annual) 



And see the result : YTM = Pre tax cost of debt = 10.924 %  


So, 


After tax cost of debt = 10.924( 1 - 0.25) = 9.193 = 8.19 % (ANSWER)

answered by: Tulsiram Garg